OpenStudy (anonymous):

MEDAL: What is the resulting ordered pair if the value of the independent variable is -1/2? f(x)=12x^2-8x

OpenStudy (noelgreco):

Replace f(x) with y, plug in 1/2 for x, and solve for y.

OpenStudy (mathstudent55):

\( f(- \dfrac{1}{2}) = 12( -\dfrac{1}{2} )^2 - 8 (- \dfrac{1}{2} ) \)

OpenStudy (radar):

@NoelGreco you tripped up on a sign. plug in -1/2.

OpenStudy (noelgreco):

oops.